OpenStudy (anonymous):

Find four consecutive integers who sum is 286

OpenStudy (anonymous):

1: x 2: x + 1 3: x + 2 4: x+ 3 4x + 6 = 286 4x = 280 x=70 plug it back in 1: 70 2: 70 +1 = 71 3: 70 +2 = 72 4: 70+ 3= 73

OpenStudy (anonymous):

sorry about taking so long i almost forgot how to do it

OpenStudy (anonymous):

Easiest way is to use symmetry, just divide by 4 and then the 4 numbers are the 2 sets of 2 integers on each side of that number.

OpenStudy (anonymous):

ok , so first, consecutive means that they are in a row. But we don't know that number so we would just put x for the first number 1:x Now, it follows each other right? So just put +1 for example if your number was 5, to find out what's next you would add 1, which is 6. so the second number that's after x, is x+1 the third number would be x + 2 the fourth number would be x + 3, do you see how now? ok, so now that we have the information above, we will combine like terms all the x's added together would be 4x. all the numbers added together would be 6. now we have 4x+6=286 let's solve like a regular algebra problem by subtracting 6 from both sides 4x=280 divide 4 x=70 We're not done yet though. This is your x, let's go back to the start where the first number is x (which we figured out was 70) let's plug in 70 and solve 1: 70 2: 70 + 1 3: 70 + 2 4: 70 + 3